OpenStudy (anonymous):

find the solution of this system of equation. -3x-7y=-66 and -10x-7y=-24

OpenStudy (lynfran):

\[-3x-7y=-66\]\[-10x-7y=-24\]by elimination method, take eq2 fram eq1

OpenStudy (lynfran):

\[7x=-42\]

OpenStudy (lynfran):

solve for x

OpenStudy (anonymous):

\[A = \left[\begin{matrix}-3 & -7 \\ -10 & -7\end{matrix}\right]\] \[B = \left(\begin{matrix}x \\ y\end{matrix}\right)\] \[C = \left(\begin{matrix}-66 \\ -24\end{matrix}\right)\] \[A ^{-1}C = B\]

OpenStudy (lynfran):

x=-42/7

OpenStudy (anonymous):

Find the inverse of matrix A and multiply it with C and you get your answer.

OpenStudy (lynfran):

then sub x=?? into the 1st eq. to find for y

OpenStudy (lynfran):

what did u get for x @Eadorno

OpenStudy (anonymous):

I got x=6 is that correct @LynFran

OpenStudy (lynfran):

it shd be -6

OpenStudy (lynfran):

x=-42/7 x=-6

OpenStudy (anonymous):

\[|A| = (-3*-7)-(-7*-10)\]\[A ^{-1} = \frac{ 1 }{ |A| }\left[\begin{matrix}-3 & -7 \\ -10 & -7\end{matrix}\right] = \left[\begin{matrix}\frac{ 1 }{ 7 } & \frac{ -1 }{ 7 } \\ \frac{ -10 }{ 49 } & \frac{ 3 }{ 49 }\end{matrix}\right]\] \[A ^{-1}C = \left[\begin{matrix}\frac{ 1 }{ 7 } & \frac{ -1 }{ 7 } \\ \frac{ -10 }{ 49 } & \frac{ 3 }{ 49 }\end{matrix}\right]\left(\begin{matrix}-66 \\ -24\end{matrix}\right) = \left(\begin{matrix}-6 \\ 12\end{matrix}\right)\]

OpenStudy (lynfran):

to find y -3(-6)-7y=-66 18-7y=-66 18+66=7y 84=7y 84/7=y 12=y i like your matrix method too @saseal

OpenStudy (anonymous):

nice thank you both @saseal @LynFran

OpenStudy (lynfran):

welcome

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

@LynFran It's my favorite method so far, works every time.